OpenStudy (anonymous):

Find the distance from the line to the given point: y=1/6x+6 point(-6,5). How could i get another point.Please help!

OpenStudy (anonymous):

Consider that point is (x,y), construct line wits those two points (-6,5) & (x,y). Because the two lines are perpendicular, product of it's gradient is -1. That's all you need :), hope this help.

OpenStudy (anonymous):

Please can you show me the steps

OpenStudy (anonymous):

Find the second line gradient first (let's use x1 and y1 for the point mentioned before): m2=(y1-5)/(x1+6) Then, from the first line, we know m1=1/6 m1*m2=-1 (1/6)*(y1-5)/(x1+6)=-1 From here we can find x1 in term of y1.